7. Assume that the differences are normally distributed. Complete parts (a) through (d) belovw Observation 1 53.4 49.4 51.7 48.1 49.1 50.6 43.954.0 54.6 48.9 54.4 53.8 49.7 51.3 46.5 53.1 (a) Determine di =Xi-Yi for each pair of data Observation 4 (Type integers or decimals.) (b) Compute d and sd (Round to three decimal places as needed.) (Round to three decimal places as needed.) (c) Test itụd 0 P-value (Round to three decimal places as needed.)
Choose the correct conclusion below. 0 A. Reject the null hypothesis. There is sufficient evidence that μd
Answer #1

Here we have data;

Observation Xi Yi d = Xi - Yi
1 53.4 54.6 -1.2
2 49.4 48.9 0.5
3 51.7 54.4 -2.7
4 48.1 53.8 -5.7
5 49.1 49.7 -0.6
6 50.6 51.3 -0.7
7 43.9 46.5 -2.6
8 54 53.1 0.9

120-S27560-2 0.9 513 = 12-124 Hypo从从 어 test sfat,A+,.CA .tt琳 (Cwnkhown d-ud Sd 2-12Y P-value - correcf opta on (A Raject heс.it.cou Maue.. Conf dence nteal: -1.513 +2-365XL 2-/24 8 -1.SI3 İ..776 The lower bound E329ユ the upper bound 「人16-26-7